Flamo:

Solve the system Equation... Attempt 2 ;-;

Flamo:

y = -x + 7 and y = 0.5(x - 3)^2

Flamo:

0.5(x - 3)^2 = -x + 7?

Hero:

Hint: Solve this for x; y = -x + 7 But it doesn't matter which way you substitute, you're going to end up with a quadratic expression either way you do it. I'd at least multiply the second equation by 2 on both sides to get rid of that decimal.

Flamo:

so Multiply 0.5 and -x by 2 @Hero?

Flamo:

I'm Lost big time.

Hero:

2(y = -x + 7) and 2(y = 0.5(x - 3)^2)

Flamo:

Oh.

Flamo:

So it's y = -2x + 14 for the first equation?

Hero:

More like 2y = -2x + 14

Flamo:

Oh.

Hero:

Anyway, Not sure how useful it is to multiply both equations by 2. Just trying to get rid of that decimal. I personally don't like decimals.

Flamo:

Oh, me too.

Flamo:

My Teacher said this to me when I asked for help, and I don't quite get it. Great, now Multiply both sides by 2 and then FOIL: -2x + 14 = x^2 - 6x + 9 Simplify: x^2 - 4x - 5 = 0 then factor to figure out the x's, then sub it x's to get y. You should end with two coordinates (x,y) where the graph's intersect.

Hero:

Yep, that looks good actually.

Flamo:

Oh, so now what?

Hero:

Well, you have to actually factor x^2 - 4x - 5

Flamo:

Oh

Flamo:

i'm so sorry, im lost right now. My Brain is Blank....

Flamo:

Rip my brian. ;-;

Hero:

Do you know how to factor quadratic expressions?

Flamo:

Uhh, Yeah... I think. I might just need a reminder.

Hero:

... a "refresher" you mean.

Flamo:

Yeah.. ;-;

Hero:

We need to factor \(x^2 - 4x - 5\) so a = 1, b = -4, and c = -5 And we need to find two numbers \(m\) and \(n\) such that \(m + n = |b|\) \(m \times n = ac\)

Flamo:

Oh, Okay..

Hero:

Basically \(ac = -5\), \(|b| = 4\) and \(x^2 - 4x - 5\) can be re-written as \(x^2 -(5 - 1)x - 5\)

Flamo:

Oh

Hero:

Then continue to factor by grouping \(x^2 - 5x + x - 5\) \(x(x - 5) + 1(x - 5)\) \((x - 5)(x + 1)\) Set each factor to zero then solve for x. Plug each x back into the original equation and solve for y.

Flamo:

Oh...

Flamo:

OH, I see

Flamo:

Wati, so Just do FOIL on (x - 5)(x - 1)?

Hero:

What I meant to do is ask, Is that the correct expression? And no we don't have to FOIL at this point. That was done already earlier.

Flamo:

Okay.

Hero:

After factoring, \(x^2 - 4x - 5\) you set each factor equal to zero, so \(x - 5 = 0 \) \(x + 1 = 0\)

Flamo:

So the Coordinates are (0, 0)?

Hero:

No, the zeroes are an application of the zero product property: \(ab = 0\). You have to solve for x for each factor.

Flamo:

Oh.

Flamo:

I'm still, So confused. Can we Start over? Step by step? because I need to Show my work. I've been on this problem for a while now.

Hero:

okay, re-post it as a new problem. I will walk you through it this time.

Flamo:

Okay..